# Heuristic Methods are Good Teachers to Distill MLPs for Graph Link Prediction

### v1

@@ Abstract
Link prediction is a crucial graph-learning task with applications including citation prediction and product recommendation. Distilling Graph Neural Networks (GNNs) teachers into Multi-Layer Perceptrons (MLPs) students has emerged as an effective approach to achieve strong performance and reducing computational cost by removing graph dependency. However, existing distillation methods only use standard GNNs and overlook alternative teachers such as specialized model for link prediction (GNN4LP) and heuristic methods (e.g., common neighbors). This paper first explores the impact of different teachers in GNN-to-MLP distillation. Surprisingly, we find that stronger teachers do not always produce stronger students: MLPs distilled from GNN4LP can underperform those distilled from simpler GNNs, while weaker heuristic methods can teach MLPs to near-GNN performance with drastically reduced training costs. Building on these insights, we propose Ensemble Heuristic-Distilled MLPs (EHDM), which eliminates graph dependencies while effectively integrating complementary signals via a gating mechanism. Experiments on ten datasets show an average 7.93% improvement over previous GNN-to-MLP approaches with 1.95-3.32$\times$ less training time, indicating EHDM is an efficient and effective link prediction method.
@@ 1 Introduction
Link prediction is a pivotal task in graph learning, with widespread usage in web applications. It involves estimating the likelihood of a link between two nodes, enabling applications including citation prediction, friend recommendation, knowledge graph completion [Schlichtkrull et al., 2018, Nathani et al., 2019, Vashishth et al., 2019], and item recommendation [Koren et al., 2009, Ying et al., 2018, He et al., 2020]. Graph neural networks (GNNs) [Kipf and Welling, 2016a, Hamilton et al., 2017, Veličković et al., 2017] have been widely used for link prediction. This includes standard generic GNNs [Kipf and Welling, 2016b, Davidson et al., 2018], as well as specifically designed GNNs for Link Prediction (GNN4LP) which are augmented with additional structural features [Zhang and Chen, 2018, Zhang et al., 2021b, Zhu et al., 2021b, Yun et al., 2021, Wang et al., 2024, Ma et al., 2024] to capture extra link-specific information [Zhang et al., 2021a]. However, all of these GNN models suffer from heavy computations in inference time due to their dependency on graphs, which require repeatedly fetching and aggregating neighboring nodes for GNN message passing [Zhang et al., 2021b].
Researchers have proposed various methods to accelerate GNN inference. There are pruning [Zhou et al., 2021, Chen et al., 2021] and quantization [Zhao et al., 2020, Tailor et al., 2020] methods that accelerate GNN inference to some extent. But the speed-up is limited because they do not resolve the graph dependency issue caused by the message passing. An alternative line of work focuses on distilling knowledge of GNN teachers into Multi-Layer Perceptron (MLP) students [Zhang et al., 2021b, Tian et al., 2022, Wu et al., 2023a, b, 2024, Lu et al., 2024, Wang et al., 2025]. By eliminating graph dependency and expensive neighborhood aggregation, MLPs offer up to 70-273$\times$ faster inference [Guo et al., 2023, Zhang et al., 2021b]. Meanwhile, GNN-distilled MLPs have been shown to achieve competitive or superior performance compared to GNN teachers in many cases.
Advances in GNN-to-MLP distillation focus primarily on node classification with only a few works extending the idea to link prediction. For example, LLP [Guo et al., 2023] distills relational knowledge from standard GNNs to MLPs to achieve better effectiveness for link prediction. Yet, the link-level GNN-to-MLP distillation remains under-explored. Besides, existing GNN-to-MLP works rely majorly on standard GNNs as teachers. In reality, state-of-the-art link prediction models often go beyond standard GNNs. GNN4LP methods, or even classical heuristic methods (e.g., common neighbors) [Kumar et al., 2020, Newman, 2001, Murase et al., 2019, Zhou et al., 2009], were shown to outperform standard GNNs in many scenarios [Li et al., 2024]. This creates an intriguing question: How does the choice of teacher model influence the resulting MLP student’s performance?
@@
In this work, we first explore how different teachers, ranging from standard GNNs to GNN4LP methods and heuristic-based approaches, affect the performance of the distilled MLP student for link prediction. Adopting the LLP framework [Guo et al., 2023] as the distillation pipeline, we have two surprising observations:
  1. 1.
    Good models might not be good teachers. Although GNN4LP models often outperform standard GNNs, the MLPs distilled from GNN4LP can actually underperform those distilled from simpler GNNs.
  2. 2.
    Heuristic methods are good teachers, even when they underperform. Classical heuristics can be effective teacher for MLPs to guide them achieve comparable performance to standard GNNs, even if the heuristic’s performance is worse than the MLP.
For the first observation, we attribute it to the limited capacity of MLPs for capturing highly intricate structural patterns that GNN4LP teachers rely on. Taking this insight along with the second observation, we propose to distill MLPs from heuristic teachers, because generating these heuristic scores is often much faster than training a GNN, which will substantially shorten overall training time as well. While heuristic teachers can effectively guide MLPs, each heuristic captures only one type of structural proximity [Mao et al., 2023], leaving certain positive links unaccounted for. One might try merging heuristics into a single “stronger” teacher, but we find that mixing multiple heuristics can overload the student and hurt its performance. Instead, we take the heuristic-to-MLP distillation one step further to propose a simple yet effective Ensemble Heuristic-Distilled MLPs (EHDM) method that trains one MLP per heuristic, then learns a gater MLP to fuse their predictions at inference time. Our gating mechanism uses only node features, thereby preserving the MLP’s fast, neighbor-free inference. This allows us to harness multiple heuristics’ complementary signals and achieve higher accuracy without reintroducing graph dependencies.
To summarize, this work makes the following contributions:
  • We show that even underperforming heuristic methods can be surprisingly good teachers to MLPs, enabling near-GNN predictive performance while reducing the cost of generating distillation guidance by an order of magnitude.
  • We propose an effective ensemble method EHDM to combine multiple heuristic-distilled MLPs via a gating mechanism which overcomes the limitations of single heuristics distillation.
  • Experiments on ten datasets (including three million-scale datasets) show an average of 7.93% performance gain over previous GNN-to-MLP methods [Guo et al., 2023], while accelerating overall training time by 1.95-3.32$\times$.
@@ 2 Preliminaries and Related Work
In this section, we discuss preliminaries and related work relevant to contextualizing our paper.
@@ 2.1 Link Prediction on Graphs
Let $G=(\mathcal{V},\mathcal{E})$ be an undirected graph with $N$ nodes and observed edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$. Node features are represented as $\mathbf{X}\in\mathbb{R}^{N\times F}$ where $\bm{x}_{i}$ is the feature vector for node $i$. Given a pair of nodes $(i,j)$, the goal of link prediction is to estimate the probability of connection, denoted as $P(i\sim j)$.
@@
The state-of-the-art solutions are based on GNNs. Early methods directly utilize the embeddings of two target nodes for prediction [Kipf and Welling, 2016b, Berg et al., 2017, Schlichtkrull et al., 2018, Ying et al., 2018, Davidson et al., 2018]. Subsequent research revealed that standard GNN methods struggle to capture link-specific information [Zhang et al., 2021a]. To address this, various approaches have been developed to enhance GNNs by incorporating additional structural information between target node pairs. Li et al. [Li et al., 2024] called these methods GNN4LP models. GNN4LP models such as SEAL [Zhang et al., 2021a], BUDDY [Chamberlain et al., 2022], NBFNet [Zhu et al., 2021a], and MPLP [Dong et al., 2023] leverage subgraph features, while Neo-GNN [Yun et al., 2021], NCN [Wang et al., 2024], and NCNC [Wang et al., 2024] utilize common neighbor information. Although these methods improve prediction accuracy, they increase the computational cost during training and inference, rendering them less suitable for large-scale web applications.
@@
Heuristic methods estimate node proximity based on structural or feature similarities [Mao et al., 2023], categorized as follows:
  • Local Structural Proximity considers 1-hop neighborhood similarity. We consider three heuristics: Common Neighbors (CN) [Newman, 2001]: $CN(i,j)=|N(i)\cap N(j)|$, Adamic-Adar (AA) [Adamic and Adar, 2003]: $AA(i,j)=\sum_{k\in N(i)\cap N(j)}\frac{1}{\log(\text{Deg}(k))}$, Resource Allocation (RA) [Zhou et al., 2009]: $RA(i,j)=\sum_{k\in N(i)\cap N(j)}\frac{1}{\text{Deg}(k)}$, where $N(i)$ is the neighbors of node $i$ and $\text{Deg}(\cdot)$ is the node degree.
  • Global Structural Proximity captures higher-order connectivity, e.g., SimRank [Jeh and Widom, 2002], Katz [Katz, 1953], and Personalized PageRank [Brin and Page, 1998]. Due to their high computational cost, we adopt Capped Shortest Path (CSP), a simplified shortest-path heuristic [Liben-Nowell and Kleinberg, 2003]: $CSP(i,j)=\frac{1}{\min(\tau,SP(i,j))}$, where $SP(i,j)$ is the shortest path length between $i$ and $j$, and $\tau$ limits the maximum distance considered [^1] With bi-directional breadth first search, we need to retrieve $\tau/2$-hops neighborhood to compute CSP..
  • Feature Proximity measures similarity in node features, assuming nodes with similar characteristics are more likely to connect [Khanam et al., 2023]. Ma et al. [Ma et al., 2024] shows MLP is good at capturing feature proximity.
@@ 2.2 GNN Acceleration
Inference acceleration has been explored through both hardware and algorithmic optimizations. On the algorithmic side, techniques such as pruning [Han et al., 2015, Zhou et al., 2021, Chen et al., 2021] and quantization [Gupta et al., 2015, Zhao et al., 2020, Tailor et al., 2020] have been widely studied to reduce model complexity and computational cost. While these techniques improve inference efficiency, they do not eliminate neighbor-fetching operation, which remains a fundamental bottleneck in GNN inference. Because GNNs rely on message passing across graph structures, even optimized models still suffer from high data dependency and irregular memory access patterns.

### v2

# Weak Models Can be Good Teachers: A Case Study on Link Prediction with MLPs

@@ Abstract
[…]
@@ 1 Introduction
Link prediction is a pivotal task in graph learning, with widespread usage in web applications. It involves estimating the likelihood of a link between two nodes, enabling applications including citation prediction, friend recommendation, knowledge graph completion [schlichtkrull2018modeling, nathani2019learning, vashishth2019composition], and item recommendation [koren2009matrix, ying2018graph, he2020lightgcn]. Graph neural networks (GNNs) [kipf2016semi, hamilton2017inductive, velivckovic2017graph] have been widely used for link prediction. This includes standard generic GNNs [kipf2016variational, davidson2018hyperspherical], as well as specifically designed GNNs for link prediction (GNN4LP) which are augmented with additional structural features [zhang2018link, zhang2021graph, zhu2021neural, yun2021neo, ncnc, linkmoe] to capture extra link-specific information [zhang2021labeling]. However, all of these GNN models suffer from heavy computations in inference time due to their dependency on graphs, which require repeatedly fetching and aggregating neighboring nodes for GNN message passing [zhang2021graph].
[…]
Advances in MLP distillation on graphs focus primarily on node classification with only a few works extending the idea to link prediction. For example, LLP [guo2023linkless] distills relational knowledge from standard GNNs to MLPs to handle link prediction. Yet, the link-level GNN-to-MLP distillation remains under-explored. Besides, existing MLP distillation works rely majorly on standard GNNs as teachers. In reality, state-of-the-art link prediction models often go beyond standard GNNs. GNN4LP methods, or even classical heuristic methods (e.g., common neighbors) [kumar2020link, newman2001clustering, murase2019structural, zhou2009predicting], were shown to outperform standard GNNs in many scenarios [heart]. This creates an intriguing question: How does the choice of teacher model influence the MLP student’s performance?
@@
In this work, we first explore how different teachers, ranging from GNNs to GNN4LP methods and heuristic-based approaches, affect the performance of the distilled MLP student for link prediction. Adopting the LLP framework [guo2023linkless] as the distillation pipeline, we have two surprising observations:
  1. Good models $\neq$ good teachers. Although GNN4LP models often outperform standard GNNs, the MLPs distilled from GNN4LP can actually underperform those distilled from simpler GNNs.
  2. Heuristic methods can be good teachers, even when they underperform. Classical heuristics can be effective teachers for MLPs to guide them achieve comparable performance to standard GNNs, even if the heuristic’s performance is worse than the MLP.
To explain these findings, we provide both empirical and theoretical analyses. The first observation arises because MLPs, with their limited capacity and lack of message passing, struggle to absorb the intricate structural patterns that GNN4LP teachers rely on. In contrast, the second observation highlights that heuristic methods, despite their simplicity and weaker standalone performance, offer complementary signals that are easier for MLPs to learn and align with.
Based on these insights, we propose to distill MLPs from heuristic teachers, which will substantially shorten overall training time. To further improve prediction accuracy, we design a simple yet effective Ensemble Heuristic-Distilled MLPs (EHDM) method that trains one MLP per heuristic, then learns a gater MLP to fuse their predictions at inference time. Our gating mechanism uses only node features, thereby preserving the MLP’s fast, neighbor-free inference. This allows us to harness multiple heuristics’ complementary signals and achieve higher accuracy without reintroducing graph dependencies. To summarize, this work makes the following contributions:
  • We show that even underperforming heuristic methods can be surprisingly good teachers to MLPs, substantially reducing the training cost while enabling near-GNN predictive performance.
  • We provide empirical and theoretical analysis to this phenomenon, revealing that a teacher is only as good as its teachable knowledge, the portion of its knowledge that can be transferred through the features accessible to the student.
  • Our work is the first to formally prove that a more expressive model class (GNN4LP) does not inherently offer better teachable knowledge than a simpler model class (GNN). This result highlights an important lesson for future GNN2MLP research: rather than focusing on teacher strength, we should aim to design models that are more teachable.
[…]
@@ 2.1 Link Prediction on Graphs
Let $G=(\mathcal{V},\mathcal{E})$ be an undirected graph with $N$ nodes and observed edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$. Node features are represented as $\mathbf{X}\in\mathbb{R}^{N\times F}$ where $\bm{x}_{i}$ is the feature vector for node $i$. Given a pair of nodes $(i,j)$, the goal of link prediction is to estimate the probability of connection, denoted as $P(Y_{ij}=1)$.
@@
The state-of-the-art solutions are based on GNNs. Early methods directly utilize the embeddings of two target nodes for prediction [kipf2016variational, berg2017graph, schlichtkrull2018modeling, ying2018graph, davidson2018hyperspherical]. Subsequent research revealed that standard GNN methods struggle to capture link-specific information [zhang2021labeling]. To overcome this limitation, GNN4LP models extend GNNs by explicitly incorporating structural information that is not inherently encoded by GNN architectures [heart]. A common example is the inclusion of global node IDs, since vanilla GNN inputs typically contain only node features. These IDs can then be exploited to compute structural signals such as the number of common neighbors [yun2021neo, ncnc] and subgraph features [zhang2021labeling, buddy, nbfnet, mplp]. Although these methods improve prediction accuracy, they increase the computational cost during training and inference, rendering them less suitable for large-scale web applications.
@@
Heuristic methods estimate node proximity based on structural or feature similarities [mao2023revisiting], categorized as follows:
  • Local Structural Proximity considers 1-hop neighborhood similarity. We consider three heuristics: Common Neighbors (CN) [newman2001clustering]: $CN(i,j)=|N(i)\cap N(j)|$, Adamic-Adar (AA) [adamic2003friends]: $AA(i,j)=\sum_{k\in N(i)\cap N(j)}\frac{1}{\log(\text{Deg}(k))}$, Resource Allocation (RA) [zhou2009predicting]: $RA(i,j)=\sum_{k\in N(i)\cap N(j)}\frac{1}{\text{Deg}(k)}$, where $N(i)$ is the neighbors of node $i$ and $\text{Deg}(\cdot)$ is the node degree.
  • Global Structural Proximity captures higher-order connectivity, e.g., SimRank [jeh2002simrank], Katz [katz1953new], and Personalized PageRank [brin1998anatomy]. Due to their high computational cost, we adopt Capped Shortest Path (CSP), a simplified shortest-path heuristic [liben2003link]: $CSP(i,j)=\frac{1}{\min(\tau,SP(i,j))}$, where $SP(i,j)$ is the shortest path length between $i$ and $j$, and $\tau$ limits the maximum distance considered [^1] With bi-directional breadth first search, we need to retrieve $\tau/2$-hops neighborhood to compute CSP..
  • Feature Proximity measures similarity in node features, assuming nodes with similar characteristics are more likely to connect [khanam2023homophily]. Ma et al. [linkmoe] shows MLP is good at capturing feature proximity.
@@ 2.2 GNN Acceleration
Inference acceleration has been explored through both hardware and algorithmic optimizations. On the algorithmic side, techniques such as pruning [han2015learning, zhou2021accelerating, chen2021unified] and quantization [gupta2015deep, zhao2020learned, tailor2020degree] have been widely studied to reduce model complexity and computational cost. While these techniques improve inference efficiency, they do not eliminate neighbor-fetching operation, which remains a fundamental bottleneck in GNN inference. Because GNNs rely on message passing across graph structures, even optimized models still suffer from high data dependency and irregular memory access patterns.